### v2

@@ 4.3 Influence of the initial $k_{0}$ on the grain reorganization mechanism
Figure 10: Distribution of the mode of rupture of the bond at the end of the test. A bond can break by chemical dissolution (the bond surface reaches a negative or a null value during the dissolution steps) or by mechanical loading (criteria presented in Equation 22 are reached).
@@
Once the two mechanisms have been described, it is important to notice in the cases of light cementations (2T or 2MB) that the final values of the $k_{0}$ are larger than the others if $k_{0}^{init}\geq k_{0}^{attr}$, see Figure 4. The reduction of the $k_{0}$ is due to the softening of the grains in this case. Considering light cementations, the softening is smaller as the Young modulus evolves only between $300$ MPa (for 2T) or $320$ MPa (for 2MB) and $80$ MPa (for uncemented material) compared to $1000$ MPa (for 13MB), for example. The Young modulus stays around the same value for light cementations and the grain reorganization can not occur. The influence of the initial value of the Young modulus is discussed in Figure 12. It seems the Young modulus has an influence, especially in the case $k_{0}^{init}\geq k_{0}^{attr}$. It appears this influence is smaller in the case $k_{0}^{init}\leq k_{0}^{attr}$.
4.4 Influence of the cementation (different percentage of contacts cemented $p_{c}$, different Young modulus $E_{m}$ and different bond sizes distribution defined by $m_{log}$ and $s_{log}$)
@@
Following the definition of the different degrees of cementation used by Sarkis et al. [24], it appears that the Young modulus $E_{m}$, the percentage of contacts cemented $p_{c}$ and the bond sizes distribution (defined by $m_{log}$ and $s_{log}$) change with the cementation. To understand the effect of the cementation on the problem those three effects must be isolated.
@@
The first aspect to be analyzed is the influence of the percentage of contacts cemented $p_{c}$. This parameter describes the ratio of contact cemented during the initial configuration (see step f in Figure 2) on the total number of contacts. Complementary simulations have been run. An 11BB sample is generated under $P_{confinement}=0.1$ MPa with an initial $k_{0}=0.2$ or $0.7$. Three different values are considered for the parameter $p_{c}=50$, $75$ or $100\%$. The results are presented in Figure 11.
@@
It appears that $p_{c}$ does not influence the evolution of $k_{0}$. A light difference can be noted for $k_{0}^{init}=0.2$ but it stays negligible compared to the difference noted with the other cementations in Figure 4. This difference can be explained by the fact the reorganization mechanisms are different depending if the initial $k_{0}$ is smaller or larger than the attractor value $k_{0}^{attr}$. As explained in Section 4.3, if $k_{0}^{init}\leq k_{0}^{attr}$ the main reorganization mechanism is the collapses of the unstable chain forces. In this case, the fact that there is less cemented contact ($p_{c}$ smaller) can weaken the global structure and affect lightly the $k_{0}$ evolution. If $k_{0}^{init}\geq k_{0}^{attr}$ the main reorganization mechanism is the grains softening. Here, the distribution of the cemented contact has no influence. To resume, it appears the parameter $p_{c}$ has no influence in the evolution of the $k_{0}$ with the debonding phenomenon.
The second parameter to be analyzed is the Young modulus $E_{m}$. To investigate the effect of this parameter, the same simulations are run but the initial value of the Young modulus can be controlled (considered the same for all cementation: $E_{m}=1$ GPa, Young modulus of the cementation 13MB). The results are shown in Figure 12.
@@
For $k_{0}^{init}\leq k_{0}^{attr}$, Figure 12a, it appears the controlled $k_{0}$ evolutions are around the $k_{0}$ evolution obtained with a 13MB cementation (not controlled). This can be understood by the fact the initial Young modulus value is the same as the one used for this cementation. It is important to note that 2T and 2MB controlled simulations give results a bit different. As explained in Section 4.3, the evolution of the $k_{0}$ is more controlled by the unstable chain forces collapses than the softening of the grains for $k_{0}^{init}\leq k_{0}^{attr}$. Lightly cemented samples are composed of fewer bonds between the grains, so the collapse of the chain forces can occur more easily. It seems the initial Young modulus has an influence on the $k_{0}$ evolution but it is not the main one.
For $k_{0}^{init}\geq k_{0}^{attr}$, Figure 12b, it appears the controlled $k_{0}$ evolutions are around the $k_{0}$ evolution obtained with a 13MB cementation (not controlled). This can be understood by the fact the initial Young modulus value is the same as the one used for this cementation. Even the 2T and 2MB controlled simulations give the same results. As explained in Section 4.3, at large initial values the evolution of the $k_{0}$ is more controlled by the softening of the grains than the unstable chain forces collapses. The grain reorganization depends strongly on the initial Young modulus of the grains in the case.
The third parameter to be analyzed is the bond size distribution controlled by $m_{log}$ and $s_{log}$ (see A). It has been emphasized in the previous discussion of this Section that the bond size distribution influences in the case $k_{0}^{init}\leq k_{0}^{attr}$, Figure 12a, and do not in the case $k_{0}^{init}\geq k_{0}^{attr}$, Figure 12b. As presented in Section 4.3, the mechanism of the $k_{0}$ evolution is the collapse of the unstable chain forces. The bond size distribution influences this phenomenon, especially the variance of the distribution $s_{log}$. Moreover, Figure 13 highlights the fact that the light cementations (2T and 2MB) dissolve totally faster than the high cementation (13BT and 13MB). This Figure represents the evolution of the $k_{0}$ with the cumulative bond surface dissolved (similar to the times). Here, the mean of the distribution $m_{log}$ controls the temporal aspect of the debonding phenomena.
@@
4.5 Influence of $P_{confinement}$
The effect of the confinement pressure $P_{confinement}$ is studied in this Section. The results are shown in Figure 14.
@@ 4.5 Influence of $P_{confinement}$
First of all, it appears that no results are obtained for the light cementations (2T and 2MB) at $P_{confinement}=10$ MPa when $k_{0}^{init}=0.2$. All of the bonds break during the initial configuration set-up as the confinement pressure is large. It is important to notice that the lateral strain $\epsilon_{lat}$ is extensive during the initial set-up for $k_{0}^{init}\leq k_{0}^{attr}$ (aims to reduce $k_{0}$ as this coefficient is around the attractor value at step g of Figure 2) and compressive for $k_{0}^{init}\geq k_{0}^{attr}$ (aims to increase $k_{0}$ as this coefficient is around the attractor value at step g of Figure 2). Because of this extensive mode and the fact that shear failure occurs more, the bonds aim to break more than in the compressive mode.
Figure 14 highlights the fact that the coordination number (the mean number of contacts per grain) increases with $P_{confinement}$. As the confining pressure is larger, the grains are more squeezed and the number of contacts increases. In DEM, this phenomenon must be dealt with carefully, especially for rigid particles. Indeed, the models used are built on the small overlap assumption. This limit is named the jammed state [42]. A new soft particle model has recently been designed to investigate deeper into this jammed state [43]. It appears here that the grain reorganization (see the evolution of the coordination number) is divided into two steps: a linear part at the beginning and a nonlinear part at the end. For a reminder, the Young modulus $E_{m}$ of the grains is decreasing with the debonding and so the sample aims to be closer to the jammed state. As the Young modulus is smaller for small cementations, this state is reached sooner. For example, let’s consider $P_{confinement}=10$ MPa (green lines) and $k_{0}^{init}=1.5$ (plain lines), the nonlinear part starts around $\xi=0.6$ for an 11BB sample and around $\xi=0.7$ for a 13MB sample. In the same idea, for $k_{0}^{init}\geq k_{0}^{attr}$ the final value of the $k_{0}$ is globally larger when $P_{confinement}$ is larger (especially for $P_{confinement}=10$ MPa, green lines). As the sample is closer to the jammed state, there is less space to reorganize. This observation is less clear in the case $k_{0}^{init}\leq k_{0}^{attr}$, as the mechanism of the $k_{0}$ evolution is dominated by the collapse of the unstable chain forces, and not the softening of the grain, see Section 4.3.
@@ 5 Discussion
5.1 The evolution of $k_{0}$ can generate failure
Figure 15 shows the evolution of $k_{0}$ between the initial and final configurations $\left(\Delta k_{0}=k_{0}^{final}-k_{0}^{initial}\right)$ for different initial $k_{0}^{initial}$ values at different confinement pressures $P_{confinement}$ and for all degrees of cementation.
@@ 5.1 The evolution of $k_{0}$ can generate failure
This evolution of the $k_{0}$ during dissolution can be very important, especially in the context of an underground reservoir. The host rock can be modeled with a simple Mohr-Coulomb criteria, illustrated in Figure 16. It appears that a reduction of $k_{0}\left(=\sigma_{II}/\sigma_{I}\right)$ aims to induce a failure as the diameter of the Mohr circle increases $\left(D_{MC}=\sigma_{I}\left(1-k_{0}\right)\right)$, where $D_{MC}$ is the diameter of the Mohr circle and $\sigma_{I}$ is the vertical stress [5]. As shown earlier in this work, this $k_{0}$ reduction appears only if $k_{0}^{init}\geq k_{0}^{attr}$.
@@ 5.2 On the Young modulus reduction assumption
As formulated by Equation 24, the Young modulus and the mechanical properties are decreased with the weathering of the rock. This assumption has a huge influence on the evolution of the $k_{0}$, especially in the case $k_{0}^{init}\geq k_{0}^{attr}$, see Section 4.2. More than a numerical investigation to understand phenomena, the assumption has a deeper meaning. Indeed, the type of contacts and the load history of the sample can be modeled by this hypothesis.
As illustrated in Figure 17, the contact can be sorted into three types: I) Frictional, II) Mixed/Cohesive (Frictional+Cemented) or III) Cemented [24]. In case I there is no cement, so the Young modulus reduction assumption is not required. Otherwise, in the first approach, the contact can be modeled as two springs (one for the grains and one for the cement) in parallel (type II) or in series (type III). In those two cases (parallel or series), the Young modulus reduction assumption should be used depending on the history of loading, as discussed in the following paragraph. The difference between the two cases is based on the final value (after the entire dissolution of the bond) of the Young modulus $E_{m}^{final}$. In case II, it will be the value assumed for frictional contact (type I) $E_{m}^{I}$. In case III, it will be null (except if a new contact type I occurs).
As explained earlier in this Section, the Young modulus reduction assumption is applied for contacts type II and III, depending on the history of loading. As illustrated in Figure 18 and coming back to the initial model with springs in parallel or in series, the cementation that occurs at pressure $P_{i}$ is considered loaded only if the current pressure applied $P_{j}$ verifies $P_{j}>P_{i}$. The Young modulus reduction hypothesis should be applied only if the bonds are assumed loaded. This discussion is summarized in Table 3. In this work, loaded contacts type II have been used. Nevertheless, some results have been obtained for non-loaded contacts type II in Figure 8.
Figure 18: Depending on the history of loading, the cementation can be loaded or not. This scheme illustrates the fact that a cementation that occurred at $P_{i}$ can be considered a) not loaded if $P_{j}=P_{i}$ or b) loaded if $P_{j}>P_{i}$, where $P_{j}$ is the current pressure applied.

### v3

@@ 4.3 Influence of the initial stress state on the main grain reorganization mechanism
Figure 14: Distribution of the mode of rupture of the bonds at the end of the test. A bond can break by chemical dissolution (the bond surface reaches a negative or a null value) or by mechanical loading (criteria presented in Equation 23 are reached).
@@
Once the two mechanisms have been described, it is crucial to note that in the cases of light cementations (2T or 2MB), the final values of the $k_{0}$ are larger than the others if $k_{0}^{init}\geq k_{0}^{attr}$, see Figure 7. In this case, the reduction of $k_{0}$ is attributable to the softening of the grains. Considering light cementations, the softening is less pronounced given that the Young modulus varies between $300$ MPa (for 2T) or $320$ MPa (for 2MB) and $80$ MPa (for uncemented material). This is, for example, in contrast to $1000$ MPa (for 13MB). In the case of light cementations, the Young modulus remains approximately constant, thereby precluding the grain reorganization. The impact of the initial Young modulus value is examined in the following Section, see Figure 16. It appears that the Young modulus exerts a notable influence, particularly in the case $k_{0}^{init}\geq k_{0}^{attr}$. This influence is less pronounced in the case $k_{0}^{init}\leq k_{0}^{attr}$.
4.4 Influence of the cementation
@@
Following the definition of the different degrees of cementation used by Sarkis et al. [53], it appears that the percentage of contacts cemented $p_{c}$, the Young modulus $E_{m}$, and the bond sizes distribution (defined by $m_{log}$ and $s_{log}$) change with the cementation. To understand the impact of the cementation on the problem these three effects must be isolated.
@@
The initial stage of the analysis focuses on the impact of the percentage of contacts that are cemented $p_{c}$. This parameter describes the ratio of contact cemented during the initial configuration (see step f in Figure 4) relative to the total number of contacts. Additional simulations were conducted for comparative purposes. An 11BB sample is generated under $P_{confinement}=0.1$ MPa with an initial $k_{0}=0.2$ or $0.7$. Three distinct values are considered for the parameter $p_{c}=50$, $75$ or $100\%$. The results are presented in Figure 15.
@@
It seems that $p_{c}$ does not affect the evolution of $k_{0}$. A slight discrepancy is observable for $k_{0}^{init}=0.2$. However, it remains negligible in comparison to the discrepancies observed with the other cementations in Figure 7.
This discrepancy can be attributed to the differing reorganization mechanisms at stake, depending on $k_{0}^{init}$. As previously outlined in Section 4.3, if $k_{0}^{init}\leq k_{0}^{attr}$ the main reorganization mechanism is the collapse of the unstable chain forces. In this case, the small amount of cemented contact ($p_{c}$ smaller) can weaken the global structure and exert a minor influence on the $k_{0}$ evolution. In the case $k_{0}^{init}\geq k_{0}^{attr}$ the main reorganization mechanism is grain softening. Here, the distribution of the cemented contact has no impact. In conclusion, it can be stated that the parameter $p_{c}$ has no influence on the evolution of the $k_{0}$ during the debonding process.
The second parameter to be analyzed is the Young modulus $E_{m}$. To investigate the effect of this parameter, simulations similar to the results presented previously are conducted, but the initial value of the Young modulus is controlled (considered the same for all cementation: $E_{m}=1$ GPa, Young modulus of the cementation 13MB). The results are presented in Figure 16.
@@
For $k_{0}^{init}\leq k_{0}^{attr}$, Figure 16a, the controlled $k_{0}$ evolutions appear to be approximately equivalent to the $k_{0}$ evolution obtained with a not controlled 13MB cementation. This can be attributed to the fact that the initial Young modulus value is identical to the one employed for this cementation. It is crucial to note that the controlled 2T and 2MB simulations give disparate results. As explained in Section 4.3, the evolution of the $k_{0}$ is predominantly influenced by the collapse of unstable chain forces for $k_{0}^{init}\leq k_{0}^{attr}$, rather than the softening of the grains. The presence of fewer bonds between the grains in lightly cemented samples facilitates the collapse of chain forces. Therefore, it can be concluded that the initial Young modulus has an impact on the evolution of $k_{0}$, although it is not the most significant factor.
For $k_{0}^{init}\geq k_{0}^{attr}$, Figure 16b, the controlled $k_{0}$ evolutions appear to be approximately equivalent to the $k_{0}$ evolution obtained with a not controlled 13MB cementation. This can be attributed to the fact that the initial Young modulus value is identical to the one employed for this cementation. Furthermore, the controlled 2T and 2MB simulations result in identical evolutions. As explained in Section 4.3, the evolution of the $k_{0}$ is predominantly influenced by the softening of the grains for $k_{0}^{init}\geq k_{0}^{attr}$, rather than the unstable chain forces collapse. The grain reorganization depends significantly on the initial Young modulus value.
The third parameter to be analyzed is the bond size distribution, which is controlled by $m_{log}$ and $s_{log}$, see Appendix A. As previously discussed in this Section, the bond size distribution affects the case $k_{0}^{init}\leq k_{0}^{attr}$, Figure 16a, but not in the case $k_{0}^{init}\geq k_{0}^{attr}$, Figure 16b. As previously presented in Section 4.3, the mechanism of the $k_{0}$ evolution can be attributed to the collapse of the unstable chain forces if $k_{0}^{init}\leq k_{0}^{attr}$. It is important to note that while the positions of the particles remain largely constant, minor fluctuations can induce changes in the chain forces. The bond size distribution exerts a notable influence on this phenomenon, particularly the variance of the distribution $s_{log}$. Furthermore, Figure 17 illustrates that the light cementations (2T and 2MB) undergo complete dissolution at a significantly faster rate than the high cementations (13BT and 13MB). This Figure represents the evolution of the $k_{0}$ with the cumulative bond surface dissolved (similar to the times). The mean of the distribution $m_{log}$ appears to control the temporal aspect of the debonding phenomena.
@@
4.5 Influence of the confinement
The impact of the confinement pressure $P_{confinement}$ is studied in this Section. The results are shown in Figure 18.
@@ 4.5 Influence of the confinement
It should first be noted that no results can be obtained for the light cementations (2T and 2MB) at $P_{confinement}=10$ MPa when $k_{0}^{init}=0.2$. The bonds are found to break during the initial configuration set-up, due to the high confinement pressure. As previously explained in Section 3, $k_{0}$ is around $k_{0}^{attr}$ at the step g of Figure 4. In order to obtain $k_{0}^{init}\leq k_{0}^{attr}$, it is necessary to extend the lateral dimension of the sample. The combination of this extension with the application of the vertical pressure at the top results in bond breakage. This phenomenon occurs less frequently in the case $k_{0}^{init}\geq k_{0}^{attr}$ (compression of the lateral dimension of the sample) because the bond breakage occurs only under shear or a tensile conditions, see Equation 23. Indeed, the sample is not under tensile loading and the shear condition occurs more frequently in extension ($k_{0}^{init}\leq k_{0}^{attr}$) than in compression ($k_{0}^{init}\geq k_{0}^{attr}$).
Figure 18 illustrates that the coordination number (the mean number of contacts per grain) increases with the confining pressure $P_{confinement}$. As the confining pressure increases, the grains are more squeezed and the number of contacts increases. In DEM, this phenomenon requires careful consideration, particularly in the case of rigid particles. Indeed, the models in question are built on the basis of the small overlap assumption. This limit is named the jammed state [82]. As illustrated in the Figure, the grain reorganization can be divided into two distinct phases: first a linear part, then a nonlinear. It should be noted that the Young modulus of the grains $E_{m}$ reduces as a result of debonding, which serves to indicate that the sample is progressing towards the jammed state. Given that the Young modulus is smaller for lighter cementations, this state is reached sooner. The examples $P_{confinement}=10$ MPa (green lines) and $k_{0}^{init}=1.5$ (plain lines) illustrate this aspect. The nonlinear portion of the curve begins to emerge at approximately $\xi=0.6$ for an 11BB sample and $\xi=0.7$ for a 13MB sample. In a similar idea, the final value of the $k_{0}$ is globally larger when $P_{confinement}$ is larger for $k_{0}^{init}\geq k_{0}^{attr}$. As the sample approaches the jammed state, the available space for reorganization is reduced. This observation is less evident in the case $k_{0}^{init}\leq k_{0}^{attr}$. Indeed, the mechanism of the $k_{0}$ evolution is dominated by the collapse of the unstable chain forces, rather than the softening of the grain, see Section 4.3.
@@ 5 Discussion
5.1 The evolution of stress state can generate failure
Figure 19 shows the evolution of $k_{0}$ $\left(\Delta k_{0}=k_{0}^{final}-k_{0}^{initial}\right)$ between the initial and final configurations for different initial $k_{0}^{initial}$ values at different confinement pressures $P_{confinement}$ and for all the degrees of cementation.
@@ 5.1 The evolution of stress state can generate failure
The evolution of the $k_{0}$ during dissolution can be of significant consequence, particularly in the context of an underground reservoir. The host rock or the triggering of slip along an existing fault can be modeled using a simple Mohr-Coulomb criterion, illustrated in Figure 20. It appears that a reduction of $k_{0}\left(=\sigma_{II}/\sigma_{I}\right)$ can result in a failure. Indeed, the diameter of the Mohr circle increases $\left(D_{MC}=\sigma_{I}\left(1-k_{0}\right)\right)$, where $D_{MC}$ is the diameter of the Mohr circle and $\sigma_{I}$ is the vertical stress and remains constant [41]. As previously demonstrated in this study, the $k_{0}$ reduction appears only in the case $k_{0}^{init}\geq k_{0}^{attr}$. Therefore, this evolution of the $k_{0}$ can potentially induce the nucleation of slip, which can lead to an earthquake if the fault is seismogenic [83].
@@ 5.2 On the Young modulus reduction assumption
As explained in Section 4.2 and illustrated in Figure 11, the assumption of a Young modulus reduction with the debonding (Equation 25) is pivotal for the evolution of the stress state of a sample subjected to dissolution of the bonds, particularly in the case of $k_{0}^{init}\geq k_{0}^{attr}$. It has been demonstrated that $k_{0}$ evolves only when the Young modulus is reduced with the debonding process. This assumption on the evolution of the Young modulus represents not only a numerical investigation of the phenomenon of dissolution but also the different types of contacts and loading history of the sample.
As illustrated in Figure 3 and detailed in Section 3, a cemented contact can be categorized into three distinct types: I) Frictional, II) Mixed/Cohesive (Frictional+Cemented), or III) Cemented [53]. For type I, there is no cement, so the Young modulus reduction assumption is not required. Otherwise, the contact can be modeled as two springs (one for the grains and one for the cement) in parallel (type II) or in series (type III). For these two types (parallel or series), the Young modulus reduction assumption should be employed, contingent on the loading history, as discussed in the following paragraph. The distinction between type II and III hinges on the ultimate value of the Young modulus $E_{m}^{final}$, post-complete dissolution of the bonds. For type II, the ultimate value will be the one which is assumed for frictional contact (type I) $E_{m}^{I}$. For type III, this value will be zero (unless a new contact of type I occurs).
As explained earlier in this Section, the Young modulus reduction assumption can be applied to Type II and III contacts, depending on the loading history. For the purposes of this discussion, it is assumed that cementation is initially generated at pressure $P_{i}$. The sample can be modeled as springs (one for the grains, one for the bonds) in parallel (Type II) or series (Type III). The cementation is considered loaded only when the applied pressure $P_{j}$ verifies $P_{j}>P_{i}$. Subsequently, the stiffness of the bond spring reduces as the dissolution process occurs. The Young modulus reduction assumption is applied to ensure the sample is representative. An illustration of this configuration is provided in Figure 7. It seems that $k_{0}$ evolves to reach the value $k_{0}^{attr}$. If $P_{j}=P_{i}$, the cementation is unloaded, and the debonding has no impact on the mechanical behavior of the sample. It would be erroneous to apply the Young modulus reduction assumption. An illustration of this configuration is provided in Figure 11. It seems that $k_{0}$ reaches the value $k_{0}^{attr}$ only in the case $k_{0}^{init}\leq k_{0}^{attr}$ and remains constant in the case $k_{0}^{init}\geq k_{0}^{attr}$. This preceding discussion is summarized in Table 3.
@@
5.3 A Young modulus dependency on the size of the bonds
As previously mentioned in Section 3, the relationship described in Equation 25 between the contact scale ($E_{m}$) and the sample scale ($\xi$) could be a matter of contention. This Section explores an alternative approach, in which the local stiffness at each contact is determined by the bond surface $E_{m}(A_{b})$ [71, 72], see Equation 31. This alternative approach eliminates the discussed relationship between the contact ($E_{m}$) and sample ($\xi$) scales.
$E_{m}(A_{b})=\frac{E_{m}^{cementation}-E_{m}^{untreated}}{A_{b,\,ref}^{cementation}}\times A_{b}+E_{m}^{untreated}$ (31)
where $E_{m}$ is the Young modulus of the contact, $E_{m}^{cementation}$ is the initial Young modulus depending on the initial degree of cementation, see Table 2, $E_{m}^{untreated}$ is the Young modulus without cementation (here $80$ MPa), $A_{b,\,ref}^{cementation}$ is a bond surface of reference depending on the initial degree of cementation, and $A_{b}$ is the bond surface of the contact. Each cementation presented in Table 1 is represented through the parameters $E_{m}^{cementation}$ and $A_{b,\,ref}^{cementation}$. This bond surface of reference $A_{b,\,ref}^{cementation}$ is computed to ensure that the mean Young modulus within the sample after the cementation, see step f) in Figure 4, is equal to $E_{m}^{cementation}$ from Table 2. Compared to the previous model, a distribution of the Young modulus appears within the sample. From the bond size distribution depicted in Appendix A and the parameters $p_{c}$, $m_{log}$, and $s_{log}$ from Table 2, the different values of $A_{b,\,ref}^{cementation}$ are available in Table 4.
Cementation 2T 2MB 11BB 13BT 13MB
$A_{b,\,ref}^{cementation}\;(\mu m^{2})$ 148 2303 4267 6577 8131
Table 4: Bond surface of reference with the different degrees of cementation.
Similar to Equation 23, the contact strength still depends on the local bond surface area $A_{b}$. If one of the rupture criteria is reached, the Young modulus of the contact is considered equal to the Young modulus without cementation ($80$ MPa). The results of the $E_{m}(\xi)$ and $E_{m}(A_{b})$ simulations are depicted in Figure 21. Similar to Figure 7, it has been shown analytically by [28] that $k_{0}$ approaches a limit given by $k_{0}^{attr}=\nu/(1-\nu)$ during weathering and stiffness reduction. Considering a window of $[0.25-0.3]$ for the value of $\nu$ [74, 75, 76], the attractor value window is $[0.33-0.43]$, in agreement with the results obtained.
a) Refer to caption b) Refer to caption
Figure 21: Evolution of $k_{0}$ with the time exposure for $P_{confinement}=0.1$ MPa for $E_{m}(\xi)$ and $E_{m}(A_{b})$ in the case $k_{0}^{init}=$ a) $0.2$ and b) $0.7$. Limits considering $k^{attr}_{0}=\nu/(1-\nu)$ are plotted with $\nu=0.25-0.3$.
Moreover, it appears that the $k_{0}$ evolution mechanisms revealed for $E_{m}(\xi)$ in Section 4.3 remain the same in the case of $E_{m}(A_{b})$. Indeed, Figure 22 emphasizes the mode of rupture of the bond at the end of the test for the distinct cementation and initial state of stress, and similarly to Figure 14, the bonds break more by mechanical loading in the case $k_{0}^{init}\leq k_{0}^{attr}$ than in case $k_{0}^{init}\geq k_{0}^{attr}$. As explained in Section 4.3, this behavior is due to the predominance of the unstable chain force in the case $k_{0}^{init}\leq k_{0}^{attr}$ and the stable chain force in the case $k_{0}^{init}\geq k_{0}^{attr}$.
Figure 22: Distribution of the mode of rupture of the bond at the end of the test. A bond can break by chemical dissolution (the bond surface reaches a negative or a null value) or by mechanical loading (criteria presented in Equation 23 are reached).
Even if the attractor configuration exists and the evolution mechanisms are similar, the final states of the stress are different in the cases $E_{m}(\xi)$ and $E_{m}(A_{b})$. In the case $k_{0}^{init}\leq k_{0}^{attr}$, the final value of $k_{0}$ appears smaller for $E_{m}(\xi)$ than for $E_{m}(A_{b})$. To understand this behavior, it is worth remembering that the particles after the bond breakage are stiffer in the case $E_{m}(\xi)$ than in the case $E_{m}(A_{b})$. While $E_{m}$ is considered equal to $80$ MPa after the breakage of the bond in the case $E_{m}(A_{b})$, the value considered in the case $E_{m}(\xi)$ is larger, see Equation 25. The rupture of the unstable chain forces being the $k_{0}$ evolution mechanism in the case $k_{0}^{init}\leq k_{0}^{attr}$, stiffer $E_{m}(\xi)$ particles located on the sides limit the collapse of the unstable chain forces compared to softer $E_{m}(A_{b})$ particles [77], reducing the amplitude of the $k_{0}$ evolution. In the same vein, the final value of $k_{0}$ appears larger for $E_{m}(A_{b})$ than for $E_{m}(\xi)$ in the case $k_{0}^{init}\geq k_{0}^{attr}$. The softening of the particles being the $k_{0}$ evolution mechanism in the case $k_{0}^{init}\geq k_{0}^{attr}$, the $k_{0}$ evolution is limited in the $E_{m}(A_{b})$ sample as $E_{m}$ is considered equal to $80$ MPa after the breakage of the bond, while the value considered in the $E_{m}(\xi)$ sample is larger, and remains decreasing, see Equation 25. The amplitude of the $k_{0}$ evolution is then smaller in the case $E_{m}(A_{b})$ than in the case $E_{m}(\xi)$.
The remaining question is to determine the accurate model between the assumptions $E_{m}(\xi)$ and $E_{m}(A_{b})$. The trivial answer is to consider the case $E_{m}(A_{b})$ as the most accurate option. Indeed, the $E_{m}(\xi)$ model implies that the contact scale ($E_{m}$) depends on the sample scale ($\xi$), which is controversial. However, coming back to the assumption of the predominance of the type II for the contacts, see Figure 3, it appears that the bond material participates in the stress transmission (not for traction), even if the contact is broken because of a mechanical loading. The Young modulus of this broken contact with residual bond material should be larger than the Young modulus considered for the uncemented grain-grain contact (type I) as the cement continues to act in parallel to the grain-grain contact and thus provide additional rigidity. The more progressive decrease of the contact modulus with $E_{m}(\xi)$ allows to take into account this effect to some extent, which is not verified by the model $E_{m}(A_{b})$. As often with models, the experimental reality is located between boundary scenarios (in this work, $E_{m}(\xi)$ and $E_{m}(A_{b})$). The existence of the attractor configuration, the mechanisms inducing $k_{0}$ evolution, and the main consequence (fault reactivation) are similar in the boundary scenarios, only the pattern of $k_{0}$ evolution is impacted by the model assumptions.